# The Effectiveness of Curvature-Based Rewiring and the Role of Hyperparameters in GNNs Revisited

### v1

@@ Abstract
Message passing is the dominant paradigm in Graph Neural Networks (GNNs). The efficiency of message passing, however, can be limited by the topology of the graph. This happens when information is lost during propagation due to being oversquashed when travelling through bottlenecks. To remedy this, recent efforts have focused on graph rewiring techniques, which disconnect the input graph originating from the data and the computational graph, on which message passing is performed. A prominent approach for this is to use discrete graph curvature measures, of which several variants have been proposed, to identify and rewire around bottlenecks, facilitating information propagation. While oversquashing has been demonstrated in synthetic datasets, in this work we reevaluate the performance gains that curvature-based rewiring brings to real-world datasets. We show that in these datasets, edges selected during the rewiring process are not in line with theoretical criteria identifying bottlenecks. This implies they do not necessarily oversquash information during message passing. Subsequently, we demonstrate that SOTA accuracies on these datasets are outliers originating from sweeps of hyperparameters—both the ones for training and dedicated ones related to the rewiring algorithm—instead of consistent performance gains. In conclusion, our analysis nuances the effectiveness of curvature-based rewiring in real-world datasets and brings a new perspective on the methods to evaluate GNN accuracy improvements.   [^1] Our code is available at https://github.com/FloTori/Revisiting-Graph-Curvature-Rewiring
@@ 1 Introduction
Machine learning on graph data, and more specifically Graph Neural Networks (GNNs), has undergone rapid development over the past few years. Both in terms of architecture variations [1, 2, 3, 4] as theoretical understanding [5, 6, 7, 8]. Due to their large flexibility GNNs have been applied in a variety of domains, from physical sciences to knowledge graphs or social sciences [9, 10]. The basis of the message passing paradigm of GNNs [11] rests on the idea of information diffusion where messages, namely the node’s feature vector, are propagated along the edges of the graph to their neighbours.
@@
This approach to GNNs has been shown to be very successful, as it combines both topological information (the graph) and specific node information (feature vectors) for predictions. However, this paradigm can also suffer from drawbacks. The receptive field of a node, i.e., the number of nodes it can receive information from, depends on the number of layers of the GNN. A low number of layers can cause under-reaching [12] where required information does not reach the nodes. On the other hand, if a node’s receptive field is too large, this can lead to oversmoothing [13, 14], where node representations are homogenised due to subsequent message passing.
@@
Recently, a lot of attention has been drawn to the problem of oversquashing [15] where structural properties of the graph called bottlenecks cause a loss of information as the messages passing through get too compressed. Research efforts have therefore focused on understanding this phenomenon [16, 17] as well as on ways to alleviate it. The most pragmatic approach consists of rewiring, i.e., a targeted addition or removal of edges, to reduce the bottlenecks in the graph.
@@
In this work, we investigate the role of graph rewiring to improve performance on node classification tasks. Discrete notions of curvatures on graphs can be used to detect the location of bottlenecks, allowing for the development of algorithmic rewiring methods such as Stochastic Discrete Ricci Flow [18]. While this point of view was presented first in [18], other definitions of curvatures such as Jost and Liu Curvature [19, 20] or Augmented Forman Curvature [21] have been proposed as alternative measures [22]. Despite that fact that work on synthetic datasets [16, 17] does indicate the occurrence of oversquashing, we here nuance these results on benchmark datasets and reevaluate the performance gains of curvature-based rewiring.
Contributions and outline
@@
The goal of this paper is to analyse the effectiveness of curvature-based rewiring to improve performances in non-synthetic graph datasets. In section 2 we introduce the relation between curvature and bottlenecks in graphs. Section 3 also experimentally verifies this condition in standard node-classification datasets, where we find that most rewired edges are not necessarily responsible for oversquashing. With this nuanced perspective on the theory underlying the rewiring algorithm, we evaluate the performance gains possible due to rewiring in section 4. When comparing performances of all curvature definitions with the performance obtained when implementing no rewiring at all, we find no measurable improvements when considering the different hyperparameter sweeps. In other words, SOTA results entirely correspond to hyperparameter tuning outliers. Our work thus calls into question the effectiveness of rewiring for (large) graph datasets and creates a starting point for further investigations on how to evaluate (GNN) accuracy improvements and how to bridge theory and experiment beyond synthetic datasets.
@@ 2.1 Preliminaries
Graphs and topology
We consider an undirected graph $G=(V,E)$ identified by a set $V$ of nodes, which are described by a feature vector $\mathbf{x}_{i}\in\mathbb{R}^{n_{0}}$ with $i\in V$, and a set of edges $E\subset V\times V$. The adjacency matrix, which describes the connection of the graph, is denoted by $A$.
@@
$\mathbf{h}_{i}^{(l+1)}=\phi^{l}\left(\mathbf{h}_{i}^{(l)},\sum_{j=1}^{n}\hat{A}_{ij}\psi^{l}\left(\mathbf{h}_{i}^{(l)},\mathbf{h}_{j}^{(l)}\right)\right)\,.$ (1)
Here, $\hat{A}$ denotes the normalised augmented adjacency matrix, i.e. the adjacency matrix $A$ is augmented by self-loops $\tilde{A}=A+I$ and then normalised by $\tilde{D}=D+I$, where $D$ denotes the diagonal degree matrix. More precisely, we have $\hat{A}=\tilde{D}^{-\frac{1}{2}}\cdot\tilde{A}\cdot\tilde{D}^{-\frac{1}{2}}$.
@@ 2.2 Discrete curvature notions on graphs
First, we consider Balanced Forman Curvature (BFc)  [18] and variations thereof:
@@
  • Modified Balanced Forman Curvature: The original code implementation provided in  [18] contained an error in the counting of four-cycles (See Appendix B for more details). We therefore implement this version as well for comparison. For an edge $i\sim j$ we have $BFc_{mod}(i,j)=0$ if $\min(d_{i},d_{j})=1$ and otherwise
@@
Secondly, we consider JLc (Jost and Liu) Curvature [20]. For an edge $i\sim j$ we have, with $s_{+}\doteq\max(s,0)$,
@@
Finally, we consider the Augmented Forman Curvature used for rewiring in [21]. Originally, the Forman curvature was introduced in [23] as a discrete analogue of the Ricci curvature that aims to mimic the Bochner–Weitzenböck decomposition of the Riemannian Laplace operator for quasiconvex cell complexes (compact regular CW complexes which are quasiconvex, i.e. the boundary of two cells can at most consist of one lower-dimensional cell). It was then adapted to graphs [24] and augmented to also include two dimensional cells such as triangles [25, 26]. The Augmented Forman curvature comes in two variants, similar to BFc.
@@
  • A variant where we also consider the four-cycle contributions to this curvature. It is important to note that, unlike the Balanced Forman curvature, the term $\square(i,j)$ in $\mathcal{AF}_{4}$—as uniquely used in [21]—counts all four-cycles located at a given edge $i\sim j$ without obstructions on diagonals inside. For an edge $i\sim j$ we have
@@
One global measure to identify potential bottlenecks in an undirected graph is given by the Cheeger constant. Small values indicate that the graph can be decomposed into two distinct sets of nodes with only a few edges between those sets. Therefore, the Cheeger constant is only non-zero for connected graphs. Due to the well-known Cheeger inequality, the Cheeger constant can be approximated by the spectral gap, i.e. the first non-zero eigenvalue of the normalised Graph Laplacian. The latter has a lower computational cost. Both notions have been linked to oversquashing. It has been shown (Theorem 6 in [18]) that diffusion-based methods [27] are limited in their ability to increase the Cheeger constant. In [16], the symmetric Jacobian obstruction between two nodes is introduced as a measure of how well information can be exchanged between those nodes. Here, higher values indicate worse information flow and it is shown that the symmetric Jacobian obstruction is bounded by the inverse of the Cheeger constant (Corollary 5.6 in [16]).
@@
There have also been attempts to rewire based on a targeted increase of the spectral gap [28]. However, global measures such as the spectral gap do not convey information about the location of local bottlenecks and thus do not necessarily help to alleviate oversquashing. Here, discrete curvature measures have, allegedly, given us a local and computable metric to estimate this effect based on the graph topology itself. In [21], a bound on connections between the neighborhoods of nodes $i$ and $j$ is derived based on $\mathcal{AF}_{4}-\mathcal{AF}_{4}^{\text{min}}$ (Proposition 3.4 in [21]). However, this bound is in general not very strict and rather heuristic. In [29], it is shown edges $i\sim j$ with an Ollivier (Ricci) Curvature $\kappa(i,j)$ close to the minimum value of $-2$ cause oversquashing (Propostion 4.4 and Theorem 4.5 in [29]). From [18] we know that $\kappa(i,j)\geq BFc(i,j)$, and through the distribution of $BFc$ in Appendix A we can see that most edges are far away from the lower limit of $-2$ of $BFc$ (and therefore also $\kappa(i,j))$. This will play an important role in the next section.
@@
The most fundamental result connecting edges with a very negative Balanced Forman curvature to oversquashing is given in [18]. Here, Theorem 4 identifies negatively curved edges as sources of distorted information for a large set of nodes in the local neighbourhood of a given edge $i\sim j$.
@@
Theorem 4 [18] Consider a MPNN as in Equation 1. Let $i\sim j$ with $d_{i}\leq d_{j}$ and assume that: 1. $|\nabla\phi_{l}|\leq\alpha$ and $|\nabla\psi_{l}|\leq\beta$ with $L\geq 2$ the depth of the MPNN. 2. There exists $\delta>0$ such that $\delta<\frac{1}{\sqrt{(d_{i}\vee d_{j})}}$ ; $\delta<\frac{1}{\gamma_{max}}$ for which $\mathrm{Ric}(i,j)\leq-2+\delta$ Then there exists $Q_{j}\subset S_{2}(i)$ satisfying $|Q_{j}|>\frac{1}{\delta}$ and for $20\leq l_{0}\leq L-$ 2 we have $\frac{1}{|Q_{j}|}\sum_{k\in Q_{j}}\left|\frac{\partial h_{k}^{(l_{0}+2)}}{\partial h_{i}^{l_{0}}}\right|<(\alpha\beta)^{2}\delta^{\frac{1}{4}}\,.$ (8)
@@
Theorem 4 above gives a reason to rewire the graph around negatively curved edges (adding triangles or four-cycles which give a positive contribution) as it increases the $\delta$ and therefore softens the bound in Eq. (8). In [18], the set $Q_{j}$ is explicitly constructed as the neighbours of $j$ that correspond to tree-like structure around the edge $i\sim j$ and the Balanced Forman curvature gives us control over the information flow to these neighbours from node $i$ along the edge $i\sim j$. The idea of the proof is to bound the left-hand side of Eq. (8) by second powers of the augmented normalised adjacency matrix $(\hat{A})^{2}_{ik}$ and then to control these contributions with knowledge about the three- and four-cycles located at edge $i\sim j$ derived from the $\delta$ bound. However, the question arises if the edges selected to be rewired during pre-processing indeed satisfy the conditions (specifically condition 2) of the theorem.
@@ 2.3 Graph curvature and rewiring
In its core idea, graph rewiring techniques encompass recent efforts to decouple the input (true) graph from the graph used for computational tasks. The original message-passing paradigm [11] for GNNs used the input graph itself for information propagation. By decoupling this input graph from the graph on which the computations are performed, one can find alleviations for the oversquashing and -smoothing problem mentioned in the previous section.
This decoupling can take many forms. One can first look at ways to modify the message passing algorithm on the original graph, for example by sampling only certain nodes to improve scalability [4], by considering multiple hops during message propagation (so replacing $A$ in Eq. 1 by powers of $A$) [30, 31, 32] or by changing the definition of a neighbour. This can be done either using shortest path distances [33] or defining importance-based neighbourhoods using random walks [34]. Alternatively, one can use a more drastic approach by changing the graph itself used for propagation, for example by pre-computing unsupervised node embeddings and using neighbourhoods defined by geometric relationships in the resulting latent space for the propagation [35], by exchanging the adjacency matrix with a sparsified version of the generalised graph diffusion matrix $\mathbf{S}$, defining a new weighted and directed graph [27], or by propagating messages along a personalised PageRank [36].
In this paper, we focus on graph-rewiring techniques as follows: Given a graph $G=(V,E)$, we construct a new graph $G^{\prime}=(V,E^{\prime})$ for performing message-passing. The set $E^{\prime}$ is constructed based on the discrete curvature measures discussed in the previous section. We treat rewiring as a pre-processing task, similar to methods like SDRF [18], BORF [29], FoSR [28], LASER [37], G-RLEF or $\nu$DReW [38].

### v2

@@ Abstract
Message passing is the dominant paradigm in Graph Neural Networks (GNNs). The efficiency of it, however, can be limited by the topology of the graph. This happens when information is lost during propagation due to being oversquashed when travelling through bottlenecks. To remedy this, recent efforts have focused on graph rewiring techniques, which disconnect the input graph originating from the data and the computational graph, on which message passing is performed. A prominent approach for this is to use discrete graph curvature measures, of which several variants have been proposed, to identify and rewire around bottlenecks, facilitating information propagation. While oversquashing has been demonstrated in synthetic datasets, in this work we reevaluate the performance gains that curvature-based rewiring brings to real-world datasets. We show that in these datasets, edges selected during the rewiring process are not in line with theoretical criteria identifying bottlenecks. This implies they do not necessarily oversquash information during message passing. Subsequently, we demonstrate that SOTA accuracies on these datasets are outliers originating from sweeps of hyperparameters—both the ones for training and dedicated ones related to the rewiring algorithm—instead of consistent performance gains. In conclusion, our analysis nuances the effectiveness of curvature-based rewiring in real-world datasets and brings a new perspective on the methods to evaluate GNN improvements.
@@ 1 Introduction
Machine learning on graph data, and more specifically Graph Neural Networks (GNNs), has undergone rapid development over the past few years. Both in terms of architecture variations (Wu et al., 2019; Kipf & Welling, 2016; Veličković et al., 2018; Hamilton et al., 2017) as theoretical understanding (Xu et al., 2018; Bronstein et al., 2021; Bodnar et al., 2021a, b). Due to their large flexibility GNNs have been applied in a variety of domains, from physical sciences to knowledge graphs or social sciences (Zhou et al., 2020; Wu et al., 2021). The basis of the message passing paradigm of GNNs (Gilmer et al., 2017) rests on the idea of information diffusion where messages, namely the node’s feature vector, are propagated along the edges of the graph to their neighbours.
@@
This approach to GNNs has been shown to be very successful, as it combines topological information (the graph) and specific node information (feature vectors) for predictions. However, this paradigm can also suffer from drawbacks. The receptive field of a node, i.e., the number of nodes it can receive information from, depends on the number of layers of the GNN. A low number of layers can cause under-reaching (Sun et al., 2022) where required information does not reach the nodes. On the other hand, if a node’s receptive field is too large, this can lead to oversmoothing (Li et al., 2018; Oono & Suzuki, 2019), where node representations are homogenised due to subsequent message passing.
@@
Recently, a lot of attention has been drawn to the problem of oversquashing (Alon & Yahav, 2021) where structural properties of the graph called bottlenecks cause a loss of information as the messages passing through get too compressed. Research efforts have therefore focused on understanding this phenomenon (Di Giovanni et al., 2023; Black et al., 2023) as well as on ways to alleviate it. The most pragmatic approach consists of rewiring, i.e., a targeted addition or removal of edges, to reduce the bottlenecks in the graph.
@@
In this work, we investigate the role of graph rewiring to improve performance on node classification tasks. Discrete notions of curvatures on graphs can be used to detect the location of bottlenecks, allowing for the development of algorithmic rewiring methods such as Stochastic Discrete Ricci Flow (Topping et al., 2021). While this point of view was presented first in (Topping et al., 2021), other definitions of curvatures such as Jost and Liu Curvature (Giraldo et al., 2023; Jost & Liu, 2014) or Augmented Forman Curvature (Fesser & Weber, 2023) have been proposed as alternative measures (Bober et al., 2023). Despite that fact that work on synthetic datasets (Di Giovanni et al., 2023; Black et al., 2023) does indicate the occurrence of oversquashing, we here nuance these results on benchmark datasets and reevaluate the performance gains of curvature-based rewiring.
Contribution
The goal of our paper is to analyse the effectiveness of curvature-based rewiring to improve performances in non-synthetic graph datasets. This is in line with Tortorella & Micheli (2023) which evaluates rewiring performances on training-free GNNs and shows that rewiring rarely confers a practical benefit for message-passing in those cases. Our work shows that this is the case in general GNNs and offers an explanation: the theoretical motivation for rewiring is not satisfied in most cases when applied to real-world datasets. Additionally, we explain the occurrence of the SOTA results when rewiring by analysing the role of hyperparameter sweeps. We find that these results can be attributed to hyperparameter tuning outliers, as has also been found in other GNN performance gains (Tönshoff et al., 2024). Our work thus calls into question the effectiveness of rewiring for graph datasets and creates a starting point for further investigations on how to evaluate (GNN) improvements and how to bridge theory and experiment beyond synthetic datasets.
Outline
@@
In section 2 we introduce the theoretical relation between curvature and bottlenecks in graphs. Section 3 experimentally verifies this condition in standard datasets, where we find that most rewired edges are not necessarily responsible for oversquashing. With this nuanced perspective on the theory underlying the rewiring algorithm, we evaluate the performance gains possible due to rewiring in section 4. When comparing performances of all curvature definitions with the performance obtained when implementing no rewiring at all, we find no consistent (at the level of distributions) improvements over the different datasets.
@@ 2.1 Preliminaries
Graphs
We consider an undirected graph $G=(\mathbb{V},\mathbb{E})$ identified by a set $\mathbb{V}$ of nodes, which are described by a feature vector $\mathbf{x}_{i}\in\mathbb{R}^{n_{0}}$ with $i\in\mathbb{V}$, and a set of edges $\mathbb{E}\subset\mathbb{V}\times\mathbb{V}$. The adjacency matrix, which describes the connection of the graph, is denoted by A.
@@
$\mathbf{h}_{i}^{(l+1)}=\phi^{l}\bigg{(}\mathbf{h}_{i}^{(l)},\sum_{j=1}^{n}\hat{\textbf{A}}_{ij}\psi^{l}\left(\mathbf{h}_{i}^{(l)},\mathbf{h}_{j}^{(l)}\right)\bigg{)}\,.$ (1)
Here, $\hat{\textbf{A}}$ denotes the normalised augmented adjacency matrix, i.e. the adjacency matrix A is augmented by self-loops $\tilde{\textbf{A}}=\textbf{A}+\textbf{I}$ and then normalised by $\tilde{\textbf{D}}=\textbf{D}+\textbf{I}$, where D denotes the diagonal degree matrix. More precisely, we have $\hat{\textbf{A}}=\tilde{\textbf{D}}^{-\frac{1}{2}}\cdot\tilde{\textbf{A}}\cdot\tilde{\textbf{D}}^{-\frac{1}{2}}$.
@@ 2.2 Discrete curvature notions on graphs
First, we consider Balanced Forman Curvature (BFc) (Topping et al., 2021) and variations thereof:
@@
  • Modified Balanced Forman Curvature: The original code implementation provided in Topping et al. (2021) contained an error in the counting of four-cycles (See Appendix B for more details). We therefore implement this version as well for comparison. For an edge $i\sim j$ we have $BFc_{mod}(i,j)=0$ if $\min(d_{i},d_{j})=1$ and otherwise
@@
Secondly, we consider JLc (Jost and Liu) Curvature(Jost & Liu, 2014). For an edge $i\sim j$ we have, with $s_{+}\doteq\max(s,0)$,
@@
Finally, we consider the Augmented Forman Curvature used for rewiring in (Fesser & Weber, 2023). Originally, the Forman curvature was introduced in (Forman, 2003) as a discrete analogue of the Ricci curvature that aims to mimic the Bochner–Weitzenböck decomposition of the Riemannian Laplace operator for quasiconvex cell complexes (compact regular CW complexes which are quasiconvex, i.e. the boundary of two cells can at most consist of one lower-dimensional cell). It was then adapted to graphs (Sreejith et al., 2016) and augmented to also include two dimensional cells such as triangles (Weber et al., 2018; Samal et al., 2018). The Augmented Forman curvature comes in two variants.
@@
  • A variant where we also consider the four-cycle contributions to this curvature. It is important to note that, unlike the Balanced Forman curvature, the term $\square(i,j)$ in $\mathcal{AF}_{4}$—as uniquely used in (Fesser & Weber, 2023)—counts all four-cycles located at a given edge $i\sim j$ without obstructions on diagonals inside. For an edge $i\sim j$ we have
@@
One global measure to identify potential bottlenecks in an undirected graph is given by the Cheeger constant. Small values indicate that the graph can be decomposed into two distinct sets of nodes with only a few edges between those sets. Therefore, the Cheeger constant is only non-zero for connected graphs. Due to the well-known Cheeger inequality, the Cheeger constant can be approximated by the spectral gap, i.e. the first non-zero eigenvalue of the normalised Graph Laplacian. The latter has a lower computational cost. Both notions have been linked to oversquashing. It has been shown (Theorem 6 in (Topping et al., 2021)) that diffusion-based methods (Gasteiger et al., 2019) are limited in their ability to increase the Cheeger constant. In (Di Giovanni et al., 2023), the symmetric Jacobian obstruction between two nodes is introduced as a measure of how well information can be exchanged between those nodes. Here, higher values indicate worse information flow and it is shown that the symmetric Jacobian obstruction is bounded by the inverse of the Cheeger constant (Corollary 5.6 in (Di Giovanni et al., 2023)).
@@
There have also been attempts to rewire based on a targeted increase of the spectral gap (Karhadkar et al., 2022). However, global measures such as the spectral gap do not convey information about the location of local bottlenecks and thus do not necessarily help to alleviate oversquashing. Here, discrete curvature measures have, allegedly, given us a local and computable metric to estimate this effect based on the graph topology itself. In (Fesser & Weber, 2023), a bound on connections between the neighborhoods of nodes $i$ and $j$ is derived based on $\mathcal{AF}_{4}-\mathcal{AF}_{4}^{\text{min}}$ (Proposition 3.4 in (Fesser & Weber, 2023)). However, this bound is in general not very strict and rather heuristic. In (Nguyen et al., 2023), it is shown edges $i\sim j$ with an Ollivier (Ricci) Curvature $\kappa(i,j)$ close to the minimum value of $-2$ cause oversquashing (Propostion 4.4 and Theorem 4.5 in (Nguyen et al., 2023)). From (Topping et al., 2021) we know that $\kappa(i,j)\geq BFc(i,j)$, and through the distribution of $BFc$ in Appendix A we can see that most edges are far away from the lower limit of $-2$ of $BFc$ (and therefore also $\kappa(i,j))$. This will play an important role in the next section.
@@
The fundamental result connecting edges with a very negative Balanced Forman curvature to oversquashing is given in Topping et al. (2021), where theorem 4 identifies negatively curved edges as sources of distorted information for a large set of nodes in the local neighbourhood of the given edge.
@@
Theorem 4 (Topping et al., 2021) Consider a MPNN as in Equation 1. Let $i\sim j$ with $d_{i}\leq d_{j}$ and assume that: 1. $|\nabla\phi_{l}|\leq\alpha$ and $|\nabla\psi_{l}|\leq\beta$ with $L\geq 2$ the depth of the MPNN. 2. There exists $\delta>0$ such that $\delta<\frac{1}{\sqrt{(d_{i}\vee d_{j})}}$ ; $\delta<\frac{1}{\gamma_{max}}$ for which $\mathrm{Ric}(i,j)\leq-2+\delta$ Then there exists $Q_{j}\subset S_{2}(i)$ satisfying $|Q_{j}|>\frac{1}{\delta}$ and for $20\leq l_{0}\leq L-$ 2 we have $\frac{1}{|Q_{j}|}\sum_{k\in Q_{j}}\left|\frac{\partial h_{k}^{(l_{0}+2)}}{\partial h_{i}^{l_{0}}}\right|<(\alpha\beta)^{2}\delta^{\frac{1}{4}}\,.$ (8)
@@
Theorem 4 above gives a reason to rewire the graph around negatively curved edges (adding triangles or four-cycles which give a positive contribution) as it increases the $\delta$ and therefore softens the bound in  equation 8. In (Topping et al., 2021), the set $Q_{j}$ is explicitly constructed as the neighbours of $j$ that correspond to tree-like structure around the edge $i\sim j$ and the Balanced Forman curvature gives us control over the information flow to these neighbours from node $i$ along the edge $i\sim j$. The idea of the proof is to bound the left-hand side of  equation 8 by second powers of the augmented normalised adjacency matrix $(\hat{\textbf{A}}_{ik})^{2}$ and then to control these contributions with knowledge about the three- and four-cycles located at edge $i\sim j$ derived from the $\delta$ bound. However, the question arises if the edges selected to be rewired during pre-processing indeed satisfy the conditions (specifically condition 2) of the theorem.
@@ 2.3 Graph curvature and rewiring
In its core idea, graph rewiring techniques encompass recent efforts to decouple the input (true) graph from the graph used for computational tasks. The original message-passing paradigm (Gilmer et al., 2017) for GNNs used the input graph itself. By decoupling this input graph from the graph on which the computations are performed, which can take many forms, one can find alleviations for the oversquashing and -smoothing problem mentioned in the previous section.
One can first look at ways to modify the message passing algorithm on the original graph, for example by sampling only certain nodes to improve scalability (Hamilton et al., 2017), by considering multiple hops during message propagation (Abu-El-Haija et al., 2019, 2020; Zhu et al., 2020) or by changing the definition of a neighbour. This can be done either using shortest path distances (Abboud et al., 2022) or defining importance-based neighbourhoods using random walks (Ying et al., 2018). A more drastic approach consists of changing the graph itself used for propagation, for example by pre-computing unsupervised node embeddings and using neighbourhoods defined by geometric relationships in the resulting latent space for the propagation (Pei et al., 2019), by exchanging the adjacency matrix with a sparsified version of the generalised graph diffusion matrix $\mathbf{S}$, by defining a new weighted and directed graph (Gasteiger et al., 2019), or by propagating messages along a personalised PageRank (Gasteiger et al., 2018).
In this paper, we focus on graph-rewiring techniques as follows: Given a graph $G=(\mathbb{V},\mathbb{E})$, we construct a new graph $G^{\prime}=(\mathbb{V},\mathbb{E}^{\prime})$ for message-passing. The set $E^{\prime}$ is constructed based on the discrete curvature measures discussed in the previous section. We treat rewiring as a pre-processing task, similar to methods like SDRF (Topping et al., 2021), BORF (Nguyen et al., 2023), FoSR (Karhadkar et al., 2022), LASER (Barbero et al., 2023), G-RLEF or $\nu$DReW (Gutteridge et al., 2023).